OpenStudy (anonymous):

Solve for x. -ax + 2b > 8 x < the quantity negative 2 times b plus 8 all over negative a x > the quantity negative 2 times b plus 8 all over negative a x < the quantity 2 times b minus 8 all over negative a x > the quantity negative 2 times b minus 8 all over a

OpenStudy (anonymous):

@welshfella , @Miss.SweetiePie

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hi

OpenStudy (misty1212):

\[-ax+2b>8\]right ?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

i will tell you what they want you to do if you tell me where you got the question from

OpenStudy (misty1212):

fair enough?

OpenStudy (misty1212):

on line class?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

here is what they want you to do it is wrong

OpenStudy (anonymous):

I need help understanding how to get the answers, I don't understand math at all.

OpenStudy (misty1212):

\[-ax+2b>8\\ -ax>8-2b\\ \]

OpenStudy (misty1212):

that is a start

OpenStudy (misty1212):

then you have to divide by \(-a\)

OpenStudy (misty1212):

so you get \[x\] on the left and \[\frac{8-2b}{-a}\] on the right, which is the same as \[\frac{2b-8}{a}\]

OpenStudy (misty1212):

you are SUPPOSED to say \[x<\frac{2b-8}{a}\] even though that answer is wrong

OpenStudy (anonymous):

oh okay, yeah you see I don't understand these questions at all.

OpenStudy (misty1212):

no wonder you are confused with this math you are being taught by idiots who do not know what a variable is

OpenStudy (anonymous):

I picked C before, but it said it was wrong.

OpenStudy (misty1212):

it is not your fault, these people at K-12 whatever that is have no business writing math questions

OpenStudy (anonymous):

Yeah, and Idk how to solve this question because of the way they wrote it.

OpenStudy (misty1212):

NONE OF THE ANSWERS ARE RIGHT

OpenStudy (anonymous):

That is what I am saying, the correct answer isn't there, even my brother said it isn't

OpenStudy (misty1212):

if you give me more information about this system you are using, i will write to them and explain that they are confused

OpenStudy (misty1212):

you cannot solve an inequality with variables like that you do not now whether to switch the inequality, because you do not know if \(-a\) is positive or negative

OpenStudy (misty1212):

write them and tell them that the problem has a mistake in it

OpenStudy (misty1212):

don't bother to try to answer it

OpenStudy (anonymous):

No worries, I will talk to my teacher about it, they have made this mistake before and my teacher saw it was. So I will let her know, and thanks!

OpenStudy (anonymous):

Can you help me with another question please?

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]